OpenStudy (anonymous):

@ikatouni Find the solution to the following equation by transforming it into a perfect square trinomial. x^2 – 12x = 13 A. {46, 55} B. {–1, 13} C. {–15, 27} D. {–21, 33}

OpenStudy (anonymous):

@aajugdar

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

x^2 - 12x = 13 x^2 - 2(6)(x) = 13 x^2 - 2(6)x + 6^2 = 13 + 6^2 (added 6^2 both sides in order to complete the square) (x-6)^2 = 13 + 36 (x-6)^2 = 49 take square root both sides x -6 = +/- 7 Thus, x - 6 = 7 or x- 6 = -7 Solve for x now. There will be 2 solutions. Good Luck! (This was a direct solution because I'm a little busy with some work. So, can't help you with discussion right now)